OpenStudy (anonymous):

What is the solution to the equation? 7(5 + 9) = 7 · x + 7 · 9

OpenStudy (anonymous):

You want to get x by itself on the right-hand side ... any term that is on the right hand-side needs to move left

OpenStudy (anonymous):

35+63=7x+63 35=7x x=35/7 x=5 If u don't get it reply.! i would explain it to u

OpenStudy (anonymous):

Thanks I understand Now :)

OpenStudy (anonymous):

can someone please help me with my question it states which is a point of the vertices for the following: y> or equal to 3x+5 y< or equal to -1/2x+2 x> or equal to -6 choice A.-6,-10 B.-6,0 C.-6,12 D.-6,-13

OpenStudy (anonymous):

If u fell my answer is good u can clck The ' Best Response ' button up there it is a reward for the answerer.